Find the limits.
f(x) = 2x2 − 3x + 15, g(x) =
3

x + 3

(a)

lim
x→3
f(x) =


(b)

lim
x→24
g(x) =


(c)

lim
x→3
g(f(x)) =
1
Expert's answer
2020-04-15T10:46:56-0400

"(a) \lim_{x\to 3}f(x) = \lim_{x\to 3} (2x^2-3x+15)\\
 \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ = \lim_{x\to 3} (2(3)^2-3(3)+15)\\
 \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ = 24\\
(b) \lim_{x\to 24}g(x) = \lim_{x\to 24} \frac{3}{x+3}\\
 \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ = \lim_{x\to 24}\frac{3}{24+3} \\
 \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ =\frac{1}{9}\\
(c) \lim_{x\to 3}g(f(x)) = \lim_{x\to 3} \frac{3}{f(x)+3}\\
 \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ = \lim_{x\to 3}\frac{3}{2x^2-3x+18} \\
 \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ =\frac{1}{9}\\"